### v1

# Differentially Private Federated Learning of Diffusion Models for Synthetic Tabular Data Generation

Abstract.
The increasing demand for privacy-preserving data analytics in finance necessitates solutions for synthetic data generation that rigorously uphold privacy standards. We introduce DP-Fed-FinDiff framework, a novel integration of Differential Privacy, Federated Learning and Denoising Diffusion Probabilistic Models designed to generate high-fidelity synthetic tabular data. This framework ensures compliance with stringent privacy regulations while maintaining data utility. We demonstrate the effectiveness of DP-Fed-FinDiff on multiple real-world financial datasets, achieving significant improvements in privacy guarantees without compromising data quality. Our empirical evaluations reveal the optimal trade-offs between privacy budgets, client configurations, and federated optimization strategies. The results affirm the potential of DP-Fed-FinDiff to enable secure data sharing and robust analytics in highly regulated domains, paving the way for further advances in federated learning and privacy-preserving data synthesis.
neural networks, diffusion models, federated learning, differential privacy, synthetic data generation, financial tabular data
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; currently under review; isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networksccs: Computing methodologies Latent variable modelsccs: Security and privacy Privacy-preserving protocolsccs: Computing methodologies Distributed computing methodologies
1. Introduction
The rapidly evolving landscape of financial regulations has amplified the significance of data analytics. Central banks and financial institutions extensively gather microdata to guide policy decisions, assess risks, and maintain global stability. However, this detailed and sensitive data introduces significant privacy challenges. Real-world tabular data, crucial for developing complex models, often contains sensitive information, necessitating compliance with stringent data protection as enforced by the California Consumer Privacy Act [^1] https://oag.ca.gov/privacy/ccpa or the European General Data Protection Regulation [^2] https://eur-lex.europa.eu/eli/reg/2016/679/oj. Despite these safeguards, there remain concerns about deploying Artificial Intelligence (AI) models due to risks of data leakage (Bender et al., 2021; Kairouz et al., 2019) and model attacks (Fredrikson et al., 2015; Salem et al., 2018), which could expose personally identifiable information or confidential training data (Bender et al., 2021; Kairouz et al., 2019; Fredrikson et al., 2015; Salem et al., 2018).
Figure 1. The Financial Diffusion (FinDiff) model (Sattarov et al., 2023) (top) and the model trained with Differential Privacy (DP) (Dwork et al., 2006) (bottom), applied for mixed-type tabular data generation.
A promising approach to mitigate these privacy risks involves generating high-quality synthetic data. Synthetic data, derived from generative processes that replicate the inherent properties of real data, can provide valuable insights while preserving privacy. Unlike conventional methods such as anonymization, synthetic data generation aims to capture the underlying patterns of real data without directly exposing sensitive information. This approach is particularly relevant in high-stakes domains like finance, where data sharing and utilization are heavily regulated.
Synthetic data facilitates compliant data sharing, promoting collaboration among researchers, domain experts, and institutions. It also alleviates usage restrictions, enabling flexible data analysis without violating confidentiality agreements or regulatory boundaries. Generating high-fidelity synthetic tabular data is essential for regulatory-compliant data sharing and modeling rare, impactful events such as fraud (Charitou et al., 2021; Barse et al., 2003) or diseases. Real-world tabular data presents specific challenges, including mixed attribute types, implicit relationships, and distribution imbalances, which necessitate advanced modeling techniques.
Figure 2. Schematic representation of the proposed DP-Fed-FinDiff model. It illustrates how each client $\omega_{i}$ independently trains a Federated Financial Diffusion (Fed-FinDiff) (Sattarov et al., 2023, 2024) model with Differential Privacy (DP) (Dwork et al., 2006). The timesteps $X_{T},\ldots,X_{1},X_{0}$ represent different stages of latent data representations in the generative reverse diffusion process. The individual model parameters $\theta_{i}^{\omega}$ are periodically aggregated on a central server to form the consolidated model $\theta^{\xi}$, which is then redistributed back to each client for the next optimization round.
Recent advancements in deep generative models have demonstrated impressive capabilities in creating diverse and realistic content across various domains, including images (Brock et al., 2018; Rombach et al., 2022), videos (Singer et al., 2022; Yu et al., 2023), audio (Wang et al., 2023), code (Li et al., 2022; Chen et al., 2022), natural language (OpenAI, 2023; Touvron et al., 2023) and tabular data (Kotelnikov et al., 2023; Sattarov et al., 2023). Notably, Denoising Diffusion Probabilistic Models (DDPMs) have shown exceptional quality and realism in synthetic image generation (Dhariwal and Nichol, 2021; Rombach et al., 2022). Training these models requires substantial computing resources and extensive data, which presents challenges when sensitive data is distributed across multiple institutions and cannot be shared due to privacy concerns. Federated Learning (FL), proposed as a solution, allows multiple devices to collaboratively train AI models under the orchestration of a central server, keeping the training data decentralized and enhancing privacy (McMahan et al., 2017; McMahan and Ramage, 2017). However, even with decentralized training, ensuring the privacy of sensitive data remains a critical issue. This is where Differential Privacy (DP) becomes essential. Differential Privacy provides a mathematical guarantee that the inclusion of a single data point in a dataset does not significantly affect the outcome of data analysis, thereby ensuring that individual data points remain confidential. By integrating DP into FL, it is possible to enhance the privacy of decentralized data. This integration ensures that the updates sent to the central server do not reveal sensitive information about individual data points.
In this work, we propose a novel learning approach that integrates: (i) Differential Privacy, (ii) Federated Learning, and (iii) Denoising Diffusion Probabilistic Models. In summary, the main contributions we present are:
  • The introduction of Differential Private Federated Financial Diffusion framework (DP-Fed-FinDiff) to create synthetic tabular data with privacy protection guarantees.
  • The framework allows for precise quantification and adjustment of the privacy budget to suit the unique confidentiality and privacy requirements of the finance industry.
  • Comprehensive empirical evaluation of DP-Fed-FinDiff using real-world financial datasets, demonstrating its effectiveness in synthesizing high-quality, privacy-compliant data.
2. Related Work
Lately, diffusion models (Cao et al., 2022; Yang et al., 2022; Croitoru et al., 2023) and federated learning (Aledhari et al., 2020; Li et al., 2021; Zhang et al., 2021a) have garnered significant research interest. The following literature review focuses on federated deep generative modeling of tabular data with differential privacy.
Deep Generative Models: Xu et al. (Xu et al., 2019) introduced CTGAN, a conditional generator for tabular data, addressing mixed data types to surpass previous models’ limitations. Building on GANs for oversampling, Engelmann and Lessmann (Engelmann and Lessmann, 2021) proposed a solution for class imbalances by integrating conditional Wasserstein GANs with auxiliary classifier loss. Jordon et al. (Jordon et al., 2018) formulated PATE-GAN to enhance data synthesis privacy, providing differential privacy guarantees by modifying the PATE framework. Torfi et al. (Torfi et al., 2022) presented a differentially private framework focusing on preserving synthetic healthcare data characteristics. Lin et al. introduced DoppelGANger, a GAN-based method for generating high-fidelity synthetic time series data (Lin et al., 2019), and later analyzed the privacy risks of GAN-generated samples, highlighting vulnerabilities to membership inference attacks (Lin et al., 2021). To handle diverse data types more efficiently, Zhao et al. (Zhao et al., 2021) developed CTAB-GAN, a conditional table GAN that efficiently addresses data imbalance and distributions. Zhang et al. (Zhang et al., 2021b) offered GANBLR for a deeper understanding of feature importance, and Noock and Guillame-Bert (Nock and Guillame-Bert, 2022) proposed a tree-based approach as an interpretable alternative. Kotelnikov et al.(Kotelnikov et al., 2022) explored tabular data modeling using multinomial diffusion models (Hoogeboom et al., 2021) and one-hot encodings, while FinDiff (Sattarov et al., 2023), foundational for our framework, uses embeddings for encoding. Recent models have emerged utilizing diffusion models to address the challenges of modeling tabular data, such as class imbalance (Schreyer et al., 2024; Roy et al., 2024) or conditional tabular data synthesis (Liu and Liu, 2024).
Federated Deep Generative Models: De Goede et al. in (de Goede, 2023) devise a federated diffusion model framework utilizing Federated Averaging (McMahan et al., 2017) and a UNet backbone algorithm to train DDPMs on the Fashion-MNIST and CelebA datasets. This approach reduces the parameter exchange during training without compromising image quality. Concurrently, Jothiraj and Mashhadi in (Jothiraj and Mashhadi, 2023) introduce Phoenix, an unconditional diffusion model that employs a UNet backbone to train DDPMs on the CIFAR-10 image database. Both studies underscore the pivotal role of federated learning techniques in advancing the domain. In the context of the mixed-type tabular data, Sattarov et al. (Sattarov et al., 2024) recently introduced FedTabDiff model that merges federated learning with diffusion models.
Differentially Private Federated Deep Generative Models: The integration of Differential Privacy (DP) proposed by Dwork et al. (Dwork et al., 2006), into federated learning (FL) frameworks has gained considerable attention, particularly in enhancing the privacy of deep generative models (Abadi et al., 2016; Dockhorn et al., 2022). Fan et al. (Fan, 2020) provide a comprehensive survey of differentially private generative adversarial networks, emphasizing their potential in FL environments. Gargary and De Cristofaro (Gargary and De Cristofaro, 2024) extend this by systematically reviewing federated generative models, including those leveraging DP. Specific implementations like Chen et al. (Chen et al., 2020)’s gradient-sanitized approach for differentially private GANs, and Lomurno et al. (Lomurno et al., 2022)’s secure data exchange framework illustrate practical applications. Meanwhile, Augenstein et al. (Augenstein et al., 2019) discuss deep generative models in FL settings to maintain privacy across decentralized datasets. Additionally, Zhang et al. (Zhang et al., [n. d.]) demonstrate the use of federated differentially private GANs in detecting COVID-19 pneumonia, showcasing a critical healthcare application. In the financial domain, initial steps have been taken for applications such as financial risk modeling (Zheng et al., 2020), fraud detection (Byrd and Polychroniadou, 2020), or anomaly detection (Schreyer et al., 2022). Recently, Sattarov et al.(Sattarov et al., 2023) introduced a financial diffusion model (FinDiff) for generating mixed-type tabular data, later extended to federated settings(Sattarov et al., 2024). Balch et al. (Balch et al., 2024) introduced a hierarchy of privacy levels for generative methods in financial applications.
These advancements highlight the synergy between differential privacy and federated learning in developing privacy-preserving deep generative models. To the best of our knowledge this is the first attempt using differentially private diffusion models in a federated learning setup for synthesizing financial mixed-type tabular data.
3. Differentially Private Federated Diffusion
This section details our proposed DP-Fed-FinDiff model, which integrates Denoising Diffusion Probabilistic Models (DDPMs) with Federated Learning (FL) and enhances it with Differential Privacy (DP) for synthetic mixed-type tabular data generation.
Gaussian Diffusion Models. The Denoising Diffusion Probabilistic Model (Sohl-Dickstein et al., 2015; Ho et al., 2020) operates as a latent variable model that incrementally perturbs data  $x_{0}\in\mathbb{R}^{d}$ with Gaussian noise  $\epsilon$ through a forward process and restores it using a reverse process. Starting from  $x_{0}$ , latent variables  $x_{1},\ldots,x_{T}$ are generated via a Markov Chain, transforming them into Gaussian noise  $x_{T}\sim\mathcal{N}(0,I)$ , defined as:
(1) $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I).$
In this context,  $\beta_{t}$ represents the noise level at timestep  $t$ Sampling  $x_{t}$ from  $x_{0}$ is expressed as  $q(x_{t}|x_{0})=\mathcal{N}(x_{t};{\textstyle\sqrt{1-\hat{\beta_{t}}}}x_{0},\hat{\beta_{t}}I)$ , where $\hat{\beta_{t}}=1-\prod_{i=0}^{t}(1-\beta_{i})$ In the reverse process, the model denoises  $x_{t}$ to recover  $x_{0}$ A neural network parameterized by  $\theta$ is trained to approximate each step as $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ , where  $\mu_{\theta}$ and  $\Sigma_{\theta}$ are the estimated mean and covariance. According to Ho et al. (Ho et al., 2020), with  $\Sigma_{\theta}$ being diagonal,  $\mu_{\theta}$ is calculated as:
(2) $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\hat{\alpha_{t}}}}\epsilon_{\theta}(x_{t},t)).$
Here,  $\alpha_{t}:=1-\beta_{t}$ $\hat{\alpha_{t}}:=\prod_{i=0}^{t}\alpha_{i}$ , and  $\epsilon_{\theta}(x_{t},t)$ represents the predicted noise component. Empirical evidence suggests that using a simplified mean squared error (MSE) loss yields better results compared to the variational lower bound  $\log p_{\theta}(x_{0})$ , as given by:
(3) $\mathcal{L}_{t}=\mathbb{E}_{x_{0},\epsilon,t}||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}.$
We employ FinDiff (Sattarov et al., 2023) as the denoising diffusion probabilistic model designed for mixed-type tabular data modality.
Federated Learning. The training of DDPMs is enhanced through Federated Learning (FL) (McMahan et al., 2017), which enables learning from data distributed across multiple clients, denoted as  $\{\omega_{i}\}^{\mathcal{C}}_{i=1}$ The overall dataset is divided into subsets,  $\mathcal{D}=\{\mathcal{D}_{i}\}^{\mathcal{C}}_{i=1}$ , each accessible by a single client  $\omega_{i}$ , with varied data distributions. We employ FedTabDiff (Sattarov et al., 2024), an extension of FinDiff (Sattarov et al., 2023), in a federated setting. A central FinDiff model  $f^{\xi}_{\theta}$ with parameters  $\theta^{\xi}$ is collaboratively learned by clients. Each client  $\omega_{i}$ retains a decentralized FinDiff model  $f^{\omega}_{\theta,i}$ and contributes to the central model’s training through synchronous updates across  $r=1,\dots,\mathcal{R}$ communication rounds. A subset of clients  $\omega_{i,r}\subseteq\{\omega_{i}\}^{\mathcal{C}}_{i=1}$ is selected each round, receiving the central model parameters  $\theta^{\xi}_{r}$ , performing $\gamma=1,\dots,\Gamma$ local optimization updates, and sending updated parameters back for aggregation. Fig. 2 illustrates the process. Federated Averaging (McMahan et al., 2017) is used to compute a weighted average of the updates, defined as:
(4) $\theta^{\xi}_{r+1}\leftarrow\frac{1}{|\mathcal{D}|}\sum_{i=1}^{\lambda}|\mathcal{D}_{i}|\;\theta^{\;\omega}_{i,r+1}\,,$
where  $\lambda$ is the number of participating clients,  $\theta^{\xi}_{r}$ the central, and  $\theta^{\omega}_{i,r}$ the client model parameters,  $r$ the communication round,  $|\mathcal{D}|$ the total sample count, and  $|\mathcal{D}_{i}|\subseteq|\mathcal{D}|$ the number of samples for client  $\omega_{i}$
Differential Privacy. The concept of Differential Privacy (DP) (Dwork et al., 2006) is a mathematical framework that ensures an algorithm’s output does not significantly change when a single data point in the input is modified, protecting individual data points from inference. Formally, a randomized algorithm $\mathcal{A}$ provides $(\varepsilon,\delta)$-differential privacy if for any two datasets $D$ and $D^{\prime}$ differing by one element, and for any subset of outputs $S\subseteq\text{Range}(\mathcal{A})$:
(5) $\mathbb{P}[\mathcal{A}(D)\in S]\leq e^{\varepsilon}\mathbb{P}[\mathcal{A}(D^{\prime})\in S]+\delta,$
where $\varepsilon$ is the privacy loss parameter (smaller $\varepsilon$ means better privacy), and $\delta$ is a small probability of failure.
Federated Learning with Differential Privacy. In the proposed DP-Fed-FinDiff model, the parameter update process is modified to incorporate the Gaussian Mechanism (Dwork et al., 2014). For each minibatch, the gradient for each individual sample $\nabla\ell(x_{t},\theta)$ is computed and then clipped individually to a maximum norm $C$. These clipped gradients are accumulated into a single gradient tensor, and Gaussian noise $\mathcal{N}(0,\sigma^{2}I)$ is added. The parameter $\sigma$ is chosen based on the desired privacy budget $\varepsilon$ and $\delta$. Each client’s local differentially private model update is computed as follows:
(6) $\theta^{\omega}_{i,r+1}=\theta^{\omega}_{i,r}-\eta(\frac{1}{|B|}\sum_{x\in B}\text{clip}(\nabla\ell(x_{t},\theta^{\omega}_{i,r}),C)+\mathcal{N}(0,\sigma^{2}I)),$
where $\eta$ denotes the learning rate and $B$ is the batch size. The central server aggregates these updates using the Federated Averaging technique defined in Equation (4), ensuring the privacy of individual client data (see Figure 2).
4. Experimental Setup
This section describes the details of the conducted experiments, encompassing used datasets, data preparation steps, model architecture including hyperparameters, and evaluation metrics.
4.1. Datasets and Data Preparation
In our experiments, we utilized the following four real-world and mixed-type tabular datasets:
  1. (1)
    Credit Default [^3] https://archive.ics.uci.edu/ml/datasets/default+of+credit+card+clients ($\mathcal{D}_{A}$): This dataset includes 30,000 customers default payments records (e.g., payment history and bill statements) from April to September 2005. Each record includes 9 categorical and 13 numerical attributes.
  2. (2)
    Census Income Data [^4] https://archive.ics.uci.edu/dataset/2/adult ($\mathcal{D}_{B}$): This dataset contains demographic information from the 1994 U.S. Census to predict whether a person earns more than \$50,000 per year. In total, there are 32,561 records each encompassing 10 categorical and 3 numerical attributes.
  3. (3)
    Philadelphia City Payments Data [^5] https://tinyurl.com/bdz2xdbx ($\mathcal{D}_{C}$): This dataset consists of 238,894 payment records. The payments were generated by 58 distinct city departments in 2017. Each payment includes 10 categorical and 1 numerical attribute(s).
  4. (4)
    Marketing Data [^6] https://tinyurl.com/zx4u8tf5 ($\mathcal{D}_{D}$) This dataset contains 45,211 customer records of a bank from 2008 to 2010, used to predict whether a client will subscribe to a term deposit. Each record includes 10 categorical and 6 numerical attributes.
Figure 3. Non-IID data distribution among (3, 5, and 10) client settings. Each bar indicates the percentage of data allocated to each client $\omega_{i}$ across four datasets.
To simulate a realistic non-IID and unbalanced data environment for federated training, each dataset is partitioned based on a categorical feature. The descriptive statistics on the non-IID data partitioning schemes of 3, 5, and 10 client settings are presented in Figure 3. To standardize the numeric attributes, we employed quantile transformations, as implemented in the scikit-learn library.  [^7] https://tinyurl.com/ht9pz8m5 For the categorical attributes, we utilized embedding techniques following the approach outlined by Sattarov et al. (2023).
4.2. Model Architecture and Hyperparameters
In the following, we detail the architecture and the specific hyperparameters chosen in DP-Fed-FinDiff model optimization.
Diffusion Model. [^8] Model parameter optimization is conducted using PyTorch v2.2.1 (Paszke et al., 2019) The architecture for all datasets consists of three layers, each comprising 1024 neurons, except for $\mathcal{D}_{C}$, which contains 2048 neurons. The models are trained for up to $R=3,000$ communication rounds utilizing a mini-batch size of 16. The Adam optimizer (Kingma and Ba, 2015) is utilized with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ The hyperparameters for the diffusion model FinDiff model are adopted from (Sattarov et al., 2023). These settings include 500 diffusion steps ( $T=500$ ) and a linear learning-rate scheduler with initial and final rates of $\beta_{start}=0.0001$ and $\beta_{end}=0.02$ , respectively. Each categorical attribute is represented as a 2-dimensional embedding.
Federated Learning. [^9] The federated learning scenario is simulated using the Flower framework v1.7.0 (Beutel et al., 2022).  [^10] https://flower.ai/docs/framework/explanation-differential-privacy.html In each communication round $r=1,\ldots,R$ , each client $\omega_{i}$ performs $\gamma=1,\dots,\Gamma$ local optimization updates on its model $\theta^{\omega}_{i}$ before sharing the updated parameters. The number of client optimization updates is evaluated across various settings $\Gamma\in[10,50,100,500,1000]$ Configurations with different numbers of clients $\lambda\in[3,5,10]$ are also examined. Four distinct federated optimization strategies are explored: Federated Averaging (FedAvg(McMahan et al., 2017), Federated Adam (FedAdam(Reddi et al., 2020), Federated Proximal (FedProx(Li et al., 2020), and Federated Yogi (FedYogi(Reddi et al., 2020). FedAvg aggregates client models by computing the weighted average of their parameters. FedAdam, an extension of the Adam optimizer, utilizes default hyperparameters $\beta_{1}=0.9$ , $\beta_{2}=0.999$ , and $\epsilon=1e-8$ FedProx introduces a proximal term to address client heterogeneity, employing a default $\mu=0.01$ Lastly, FedYogi adapts the Yogi optimizer using default parameters $\beta_{1}=0.9$ , $\beta_{2}=0.999$ , and $\epsilon=1e-8$
Differential Privacy. [^11] For training and accounting of the differential privacy we use Opacus v1.4.1 (Yousefpour et al., 2021). We adopt the privacy settings from (Dockhorn et al., 2022), training models with $\varepsilon$ values of $\{0.2,1,10\}$, corresponding to high, moderate, and low privacy levels. The probability of information leakage is set to the reciprocal of the number of training samples, $\delta=N^{-1}$, a common heuristic in practice.
(a) Non-IID Results
(b) IID Results
Figure 4. Comparative evaluation of local optimization updates $\Gamma\in[10,50,100,500,1000]$ between (a) non-IID and (b) IID settings, highlighting their impact on privacy, utility, and fidelity. In IID settings, increasing the number of local updates consistently improves performance, whereas, in non-IID settings, performance degrades after reaching a certain threshold.
Figure 5. Heatmaps illustrating the impact of Differential Privacy (DP) budgets $\varepsilon\in[0.2,1,10]$ and $\infty$ (no DP) on various datasets across three key metrics: Privacy, Utility, and Fidelity in non-IID settings. It is observed that as the DP budget decreases (i.e., as $\varepsilon$ values lower from $\infty$ to 0.2), privacy protection improves, while fidelity and utility decline.
(a) Original
(b) Synthetic ($\varepsilon$$\infty$)
(c) Synthetic ($\varepsilon$=10)
(d) Synthetic ($\varepsilon$=1)
(e) Synthetic ($\varepsilon$=0.2)
Figure 6. t-SNE visualization of (a) original Credit Default data ($\mathcal{D}_{A}$), (b) synthetic data generated without differential privacy (DP), and (c, d, e) synthetic data generated with DP using federated optimization across 5 clients ($\lambda$=5). As privacy levels increase (reflected by a decrease in $\varepsilon$), the structural integrity of the synthetic data diminishes, resulting from the increased DP noise.
4.3. Evaluation Metrics
A comprehensive set of standard evaluation metrics, including privacy, utility, and fidelity, is employed to evaluate the model’s effectiveness. These metrics represent diverse aspects of data generation quality, providing a holistic view of model performance.
Privacy. [^12] The estimation of privacy risks was conducted using the Anonymeter library (Giomi et al., 2023). The privacy metric quantifies the extent to which synthetic data prevents the identification of original data entries. In this study, privacy is assessed using three key indicators of factual anonymization as outlined by the GDPR (Article 29 Data Protection Working Party, 2014). Specifically, we employ privacy evaluators to measure the risks of (i) singling out, (ii) linkability, and (iii) inference that could potentially affect data donors following the release of synthetic datasets. These risk measurements are defined as follows, where $s$ denotes the synthetic dataset and $x$ represents the real dataset:
  • Singling Out Risk quantifies the probability $SOR(x,s)$ that a synthetic record uniquely corresponds to a real record. This risk assesses the likelihood that an individual in the real data can be identified based on a unique synthetic entry. Results include a 95% confidence interval.
  • Linkability Risk measures the proportion of successful attribute linkages $LR(x,s)$ between synthetic and real records. This risk evaluates the potential for linking a synthetic record to a real one by matching shared attributes, using a subset of 10 attributes (6 for dataset $\mathcal{D_{C}}$).
  • Inference Risk evaluates an attacker’s ability to predict a secret attribute using auxiliary data, quantified by model accuracy $IR(x,s)$. This risk measures how well an adversary can infer unknown information, with each column as a secret and others as auxiliary data.
For each evaluator, the risk is estimated by performing 500 attacks on each record, and the overall privacy risk is computed as the mean score across all synthetic data points. The comprehensive privacy score is the aggregated risk from all three evaluators:
(7) $\Pi=\frac{1}{3}(SOR(x,s)+LR(x,s)+IR(x,s)).$
This empirical, attack-based evaluation framework ensures a robust assessment of privacy in synthetic data, reflecting real-world privacy risks more accurately than traditional metrics.
Utility. The effectiveness of synthetic data is determined by its utility, a measure of how functionally equivalent it is to real-world data. This utility is quantified by training machine learning models on synthetic datasets and then assessing their performance on original datasets. In this study, we operationalize utility as the performance of classifiers trained on synthetic data ($S_{Train}$) that shares dimensional consistency with the real training set and then evaluated against the actual test set ($X_{Test}$). This process evaluates the synthetic data’s efficacy in replicating the statistical properties necessary for accurate model training. The average accuracy across all classifiers is computed to represent the overall utility of the synthetic data, formalized as:
(8) $\Phi=\frac{1}{N}\sum_{i=1}^{N}\Theta_{i}(S_{Train},X_{Test}).$
Here, $\Phi$ represents the utility score, and $\Theta_{i}$ denotes the accuracy of the $i$-th classifier. To provide a comprehensive evaluation, we selected $N$=5 classifiers for this study, namely Random Forest, Decision Trees, Logistic Regression, Ada Boost, and MLP Classifier.
Fidelity. [^13] The row fidelity computation was performed using Dython library v0.7.5 (Zychlinski, 2018). Fidelity assesses how closely synthetic data emulates real data, considering both column-level and row-level comparisons. For column fidelity, the similarity between corresponding columns in synthetic and real datasets is evaluated. Numeric attributes employ the Wasserstein similarity, represented as $WS(x^{d},s^{d})$ , to measure the distance between the distributions of numeric attributes. The Jensen-Shannon divergence, denoted as $JS(x^{d},s^{d})$ quantifies differences in categorical attributes. These metrics were combined to form the column fidelity score $\Omega_{col}$ as:
(9) $\Omega_{col}=\begin{cases}1-WS(x^{d},s^{d})&\text{if $d$ is num.}\\ 1-JS(x^{d},s^{d})&\text{if $d$ is cat.}\end{cases}$
The overall fidelity for columns in synthetic dataset $S$ is the mean of $\Omega_{col}(x^{d},s^{d})$ across all attributes. Row fidelity focuses on correlations between column pairs. For numeric attributes, the Pearson Correlation between pairs $\rho(x^{a},x^{b})$ is used. The discrepancy in correlations for real and synthetic pairs, $PC(x^{a,b},s^{a,b})=|\rho(x^{a},x^{b})-\rho(s^{a},s^{b})|$ , quantifies this aspect. The Theil U coefficient, also known as Theil’s uncertainty coefficient, similarly quantifies the association between two categorical variables, denoted as $TU(x^{a,b},s^{a,b}$) . Similarly to the column fidelity, these metrics are combined to form the row fidelity:
(10) $\Omega_{row}=\begin{cases}1-PC(x^{a,b},s^{a,b})&\text{if $d$ is num.}\\ 1-TU(x^{a,b},s^{a,b})&\text{if $d$ is cat.}\end{cases}$
The total row fidelity for the dataset $S$ is the average of $\Omega_{row}(x^{a,b},s^{a,b})$ across all attribute pairs. Finally, the aggregate fidelity score, denoted as $\Omega(X,S)$ , is the mean of column and row fidelity.
5. Experimental results.
This section presents the results of the experiments, demonstrating the efficacy of the DP-Fed-FinDiff model and providing quantitative analyses. The conducted experiments are accompanied by three Research Questions (RQ) described below.
Figure 7. Singling Out, Linkability, and Inference risk evaluation under various privacy budgets ($\varepsilon$$\infty$ denotes no DP) across all datasets ($\mathcal{D}_{A}$-$\mathcal{D}_{D}$). It is observed that the privacy risk metrics increase as the privacy budget increases.
RQ 1: How does varying the number of local optimization updates impact the training of diffusion models for tabular data in a federated learning setup employing differential privacy?
Minimizing model exchange frequency is crucial for privacy while maintaining the global model’s generalization capabilities. We define $\Gamma$ as the local updates performed by a client $\omega_{i}$ at each round $r$ before synchronization. We selected $\Gamma$ values from $[10,50,100,500,1000]$ and conducted experiments in both IID and non-IID settings, fixing the privacy budget at $\varepsilon=1$ with five federated clients ($\lambda=5$).
Results: We observed an increase in the average privacy risk across all datasets with more local optimization updates in both IID and non-IID settings, peaking at 500 updates (see Figure 4). The results suggest that more local updates reduce privacy protection.
For utility and fidelity, increasing the number of client updates in the IID setting (see 4(b)) consistently improved performance. Each client received an IID data partition, which helped maintain model quality. However, in the non-IID setting (see 4(a)), performance declined after a threshold of $\Gamma=100$. The decline is caused by heterogeneous data distributions among clients. As a result, the clients drift away from a globally optimal model (Karimireddy et al., 2019), leading to unstable and slow convergence.
Our findings indicate that 100 local updates provide an optimal balance between privacy, utility, and fidelity in both IID and non-IID settings. Additionally, this choice significantly reduced the overall training time from 28 hours ($\Gamma=1000$) to 3.5 hours ($\Gamma=100$).
RQ 2: How does the application of Differential Privacy affect the generation of mixed-type tabular data and the associated privacy risks in a federated learning setting?
We assessed the impact of differential privacy on fidelity, utility, and privacy risks across three privacy budgets $\varepsilon\in[0.2,1,10]$ and without DP ($\varepsilon=\infty$) in a non-IID setting. The number of federated clients was fixed at $\lambda=5$ with client optimization rounds set to $\Gamma=100$. Privacy risks were measured by estimating Singling Out, Linkability, and Inference risks.
Figure 8. Privacy, utility, and fidelity evaluation using a varying number of federated clients $\lambda\in[3,5,10]$ across all datasets ($\mathcal{D}_{A}$-$\mathcal{D}_{D}$) and non-IID setting. We kept the privacy budget $\epsilon=1$, local updates $\Gamma=100$ and FedAvg strategy fixed.
Figure 9. Privacy, utility, and fidelity evaluation using different federated strategies across all datasets ($\mathcal{D}_{A}$-$\mathcal{D}_{D}$) and non-IID setting. We kept the privacy budget $\epsilon=1$, local optimization updates $\Gamma=100$ and number of clients $\lambda=5$ fixed.
Results: Smaller privacy budgets (lower $\varepsilon$ values) enhance privacy protection but degrade data quality across all datasets (see Figure 5). Notably, a moderate privacy budget ($\varepsilon=1$) increases privacy by 34% while reducing utility by 15% and fidelity by 14% compared to the non-DP scenario ($\varepsilon=\infty$).
Our qualitative analyses, presented in Figure 6, support these observations. The 2D t-SNE representations demonstrate the preservation of sample relationships. Without DP (see 6(b)), sample clusters closely resemble those in the original data (see 6(a)). Introducing DP gradually alters the data structure, with more pronounced changes as privacy levels increase (see 6(c) to 6(e)). This progression reflects the trade-off between data utility and privacy enhancement.
Additionally, lower DP budgets ($\varepsilon=0.2$ and $\varepsilon=1.0$) mitigate Singling Out, Linkability, and Inference risks, demonstrating the need for stringent DP constraints to minimize re-identification, linkage, and inference threats (see Figure 7). Higher DP budgets ($\varepsilon=10$ and $\infty$) reduce noise, enhancing data utility but compromising privacy. These findings highlight the trade-off between privacy and utility in synthetic data generation, emphasizing the importance of careful DP parameter selection.
RQ 3: What is the impact of varying the number of federated clients and different strategies on the quality of generated mixed-type tabular data with differential privacy?
We evaluated the impact on fidelity, utility, and privacy of synthetic data with three settings of federated clients $\omega\in[3,5,10]$ using non-IID data partitions. The strategies compared included FedAvg, FedAdam, FedProx, and FedYogi. The number of client optimization rounds was fixed at $\Gamma=100$ with a privacy budget of $\varepsilon=1$.
Results: Increasing the number of federated clients enhances privacy protection, as it becomes more difficult to infer individual data points with more clients involved (see Figure 9).
Additionally, fidelity and utility scores also show improvement with a higher number of clients. We attribute this to the regularizing effect of differential privacy, which reduces the client drift effect and results in more consistent local models. However, these benefits have limits. When expanding from five to ten clients, the gains in utility become marginal and can even decline. This performance degradation is likely due to the increased complexity of aggregating updates from a larger number of clients.
In terms of optimization strategies, there is no clear best choice when considering the trade-offs between fidelity, utility, and privacy. The strategies perform nearly identically in fidelity and utility, with slight decreases in performance due to unbalanced data distribution across clients (see Figure 9). However, significant differences are observed in privacy performance across different datasets, indicating that dataset characteristics significantly influence the effectiveness of these federated learning strategies.
In summary, the evaluation of RQ 1, 2, and 3 reveals the trade-offs in federated learning with differential privacy, highlighting the critical balance required between privacy and data quality, as well as the influence of optimization strategies and federated configurations on the overall DP-Fed-FinDiff model performance.
6. Conclusion
This study introduces the DP-Fed-FinDiff framework, a novel integration of (i) Differential Privacy (DP), (ii) Federated Learning (FL), and (iii) Denoising Diffusion Probabilistic Models (DDPMs) for generating high-fidelity synthetic tabular data. The framework addresses the critical need for privacy-preserving data generation in finance and other sensitive but high-stake domains.
Our comprehensive evaluations revealed the trade-offs between data quality and privacy. The model proves to be a robust solution for generating privacy-preserving synthetic data, paving the way for secure data sharing and advanced analytics in high-stakes environments. Future work will explore adaptive strategies to dynamically balance privacy and data quality, further enhancing the applicability of federated learning in diverse settings.